If a 10 percent rise in price leads to a reduction in quantity demanded of more than 10 percent,

demand is inelastic.

demand is elastic.

elasticity of demand is unitary.

None of the above is correct.

If the price of tuna fish increases from 50 cents to 60 cents per can and the quantity demanded decreases from 100 cans to 50 cans, then the tuna fish producer could increase its total revenue by

lowering price.

leaving price the same.

raising price.

decreasing quantity supplied.

decreasing supply.

When the price of running shoes goes from $100 to $80, the quantity demanded increases from 20 to 30 million. Over this price range, the absolute value of the price elasticity of demand is

1.80.

1.

0.55.

1.25.

2.50.

When economists say the price elasticity of supply is elastic, they mean that if prices rise

the change in quantity supplied is relatively small compared to the change in price.

suppliers are willing to produce much larger amounts of their good.

suppliers are willing to produce only a small amount more of their good.

consumers are willing to purchase much larger quantities of the good.

All things equal, the price elasticity of supply

is the same for the short run and the long run.

will be greater in the short run than in the long run.

approaches zero in the long run.

will be greater in the long run than in the short run.
Answer #1

1) QD changes by more than 10% so the demand is elastic

option(B)

2) E = [50-100/(50+100/2)] / [60-50/(60+50/2)] = [50/75] / [10/55] = 0.6666/0.1818 = 3.6

The demand is elastic so the total revenue can be increased by lowering price

option(A)

3) E = [30-20/(30+20/2)] / [80-100/(80+100/2)] = [10/25] / [20/90] = [0.4/0.2222] = 1.8

option(A)

4) suppliers are willing to produce much larger amounts of their good

option(B)